OpenStudy (anonymous):

Find the vertex of f(x)= -x^2-6x

OpenStudy (anonymous):

\[-\frac{b}{2a}\] is the first coordinate

OpenStudy (anonymous):

\[-\frac{-6}{2\times (-1)}=-3\]

OpenStudy (anonymous):

Does this go for everytime I want to find the Vertex?

sam (.sam.):

yes, you can use completing the square too